OpenStudy (anonymous):

Solve the exponential equation, and express in exact form only. e^(4x^2 - 11) = 8

OpenStudy (amistre64):

how do we undo that "e" to get to the "x" ? ahh ... thats better :)

OpenStudy (anonymous):

take the ln of both sides

OpenStudy (anonymous):

I've been trying to figure that out? I know that if they have the same base I would be able to solve it, but I'm quite confused.

OpenStudy (amistre64):

dapauls right, we can take the natural log of both sides. the ln function undoes the "e" function

OpenStudy (anonymous):

In(e^4x^2-11)/In8

OpenStudy (amistre64):

\[exp(4x^2 - 11) = 8\] \[ln(exp(4x^2 - 11)) = ln(8)\] \[4x^2 - 11 = ln(8)\] and the rest is what it is

OpenStudy (anonymous):

exactly amistre64

OpenStudy (anonymous):

\[4x^2 - 11= 2.7094 \] I would add 11 and divide by 4, then take the square.

OpenStudy (amistre64):

yep, and depending on how exacting your text requires; it might want a plus/minus applied to the sqrt

OpenStudy (amistre64):

x^2 = 7 x = sqrt(7) is called the ... principal root i think, but: x = +- sqrt(7) gives all the results

OpenStudy (anonymous):

ok. Thanks!